Fatima is having a party and needs the following:


· 27 chocolates


· 25 Packets of chips




When answering the following questions, always write the ratios in simplest form.



3.1 What is the ratio of the number of chocolates Fatima buys to the number of chocolates she gets for free?( 1 )



3.2 What is the ratio of the number of packets of chips Fatima buys to the number of packets she gets for free?( 2 )



3.3 How many chocolates should Fatima buy so that she has 27 chocolates in the end?( 1 )



3.4 How many packets of chips should Fatima buy so that she ends up with 25 packets?( 1 )



3.5 What is the total price Fatima pays for all her goods?( 3 )






1
Expert's answer
2021-11-04T16:42:04-0400

3.1

"2:1"

3.2


"4:1"


3.3


"\dfrac{27}{2+1}\cdot2=18"

 Fatima should buy18 chocolates.


3.4


"\dfrac{25}{4+1}\cdot4=20"

 Fatima should buy 20 packets of chips.


3.5


"18(8.60)+20(5.80)=270.8"

Fatima pays R270.8 for all her goods.